OpenStudy (yellowlegoguy99):

Which expression is equivalent to (7^3)^-2? 1/7 x 7 x 7 x 7 x 7 x 7 7 1/6 x 6 x 6 x 6 x 6 x 6 x 6 1/7

OpenStudy (yellowlegoguy99):

@SolomonZelman @AnyTipzWouldHelp

OpenStudy (solomonzelman):

\(\Huge\color{red}{ \rm (7^3)^{-2}=\frac{1 }{(7^3)^{2}} =\frac{1 }{7^{(3 \times 3)}} =? }\)

OpenStudy (anonymous):

Dangit solomon, you type to fast. lol

OpenStudy (yellowlegoguy99):

Is it A?

OpenStudy (anonymous):

I'll let him explain it, i'm on a couple other question right now, i can't sorry.

OpenStudy (shiraz14):

@SolomonZelman : That's incorrect - (7^3)^2 ≠ 7^(3x3).

OpenStudy (solomonzelman):

it is none of your options, unfortunately.

OpenStudy (shiraz14):

Hint: (a^b)^c = a^(bc)

OpenStudy (yellowlegoguy99):

o: 2 Math Geeks fighting!! A NEVER BEFORE SEEN MOMENT!!!!!

OpenStudy (yellowlegoguy99):

Isn't it 7^(3x2)?

OpenStudy (solomonzelman):

shiraz, you are confusing \[\Huge a^{b^c}≠a^{b+c}\]

OpenStudy (solomonzelman):

I am relevantly bad at math, but this I do know .

OpenStudy (solomonzelman):

My bad

OpenStudy (solomonzelman):

it's 3 times 2...

OpenStudy (solomonzelman):

the exponent I mean... ; sorry.

OpenStudy (yellowlegoguy99):

Okay, then it's A. Thanks

OpenStudy (shiraz14):

@SolomonZelman : Yes, your last post is correct. Indices work differently from normal mathematical operators. Some laws which you need to know are as follows: 1. a^b * a^c = a^(b+c) 2. (a^b)^c = a^(bc) 3. a^(b^c) = (a^b)^c

OpenStudy (anonymous):

also @yellowlegoguy99 if your ever stuck in math, it's valuable to learn back roads to getting answers. If you have a calculator, just solve (7^3)^-2 on your calculator, then for each possible answer, work them out, so for A for instance, multiply 7 all those times, then divide by 1... you get the same answer, and therefore, you can induce that A is the right answer.

OpenStudy (anonymous):

That's what I did lol

OpenStudy (solomonzelman):

yes, shiraz I know the laws, I just really messed up on this one. dky .

OpenStudy (anonymous):

It happens to everyone, no sweat.

OpenStudy (shiraz14):

@SolomonZelman : No worries - we all make mistakes at times - take care.

OpenStudy (solomonzelman):

:)

OpenStudy (shiraz14):

My typo for Law #3 - I meant to type: a^b^c = a^(b^c) Sorry for the typo.